# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ 3.2 Diffusion-based categorical reparameterization
Given that ReinMax (liu2023reinmax) provides significant improvements over the hard ST estimator, we also derive a ReinMax version of our diffusion-based reparameterization trick. First, we may re-interpret our algorithm as a composition of the reparameterization trick for continuous distribution composed with the Straight-Through gradient trick. Indeed, by the marginalization property, $\pi_{\!\theta}(x_{0})=\int\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})\,\pi^{\theta}_{{t_{1}}}(x_{t_{1}})\mathrm{d}x_{t_{1}}$, and we can write using the tower property that $\mathbb{E}_{\pi_{\!\theta}}[f(X_{0})]=\mathbb{E}_{\pi^{\theta}_{{t_{1}}}}[h_{\theta}(X_{t_{1}})]$ where $h_{\theta}(x_{t_{1}})\coloneqq\sum_{x_{0}}f(x_{0})\,\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})$. The Gaussian mixture $\pi^{\theta}_{{t_{1}}}$ can be reparameterized approximately using the map $T^{\theta}_{t_{1}}$ and thus we can approximate the gradient $\nabla_{\!{\theta}}\mathbb{E}_{\pi^{\theta}_{{t_{1}}}}[h_{\theta}(X_{t_{1}})]_{|\theta^{\prime}}$ with
$\nabla_{\!{\theta}}h_{\theta}(T^{\theta^{\prime}}_{{t_{1}}}(X_{t_{1}}))_{|\theta^{\prime}}+\operatorname{J}_{\theta}\!T^{\theta}_{t_{1}}(X_{1})^{\top}_{|\theta^{\prime}}\nabla_{\!{x}}h_{\theta^{\prime}}(T^{\theta^{\prime}}_{t_{1}}(X_{1}))\;.$
@@
The only intractable terms are the gradient w.r.t. $\theta$ and $x$ of the conditional expectation $h_{\theta}$. These gradients, by abstracting away $T^{\theta^{\prime}}_{t_{1}}(X_{t_{1}})$ which is not differentiated through, are $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})$ and $\nabla_{\!{x}}h_{\theta}(x_{t_{1}})$ and are a specific case of differentiating an expectation w.r.t. the parameters of a categorical distribution, which in this case is $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})$. Here by using the Straight-Through approximation (4) we recover our hard gradient estimator (15); i.e. $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})\approx\nabla_{\!{\theta}}f(\hat{x}^{\theta}_{0}(x_{t_{1}},{t_{1}}))$ and $\nabla_{\!{x}}h_{\theta}(x_{t_{1}})\approx\nabla_{\!{x_{t_{1}}}}f(\hat{x}^{\theta}_{0}(x_{t_{1}},{t_{1}}))$. Our ReinMax-based estimator thus consists in using ReinMax (5) instead of hard ST as estimator for $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})$. We coin this gradient estimator ReDGE-Max. When using a single diffusion step, i.e. ${t_{1}}=1$, $h_{\theta}$ is constant and equal to $\mathbb{E}_{\pi_{\!\theta}}[X]$ because of the boundary condition $\alpha_{1}=0$ and we recover ReinMax as a special case.
3.4 Parameter dependent $\pi_{1}$.
In the previous construction, the terminal distribution $p_{1}$ is fixed to a standard Gaussian $\pi_{1}=\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$. In our setting, however, we can exploit the factorization (2) to choose a parameter–dependent Gaussian distribution $\pi^{\theta}_{1}$ that best approximates $\pi_{\!\theta}$ in the maximum–likelihood sense. Formally, we take $\pi^{\theta}_{1}$ with factorized density $\pi^{\theta}_{1}(x)=\prod_{i=1}^{L}\mathrm{N}(x^{i};\mu^{i}_{\theta},\mathrm{Diag}(v^{i}_{\theta}))$ where for all $i\in[L]$, $(\mu^{i}_{\theta},v^{i}_{\theta})\in{\mathbb{R}}^{K}\times{\mathbb{R}}^{K}_{>0}$ and $\mathrm{Diag}(v^{i}_{\theta})\in{\mathbb{R}}^{K\times K}$ is a diagonal matrix with $v^{i}_{\theta}$ as diagonal entries. The parameters are then defined as any solution of the maximum–likelihood problem
$\{(\mu^{i}_{\theta},v^{i}_{\theta})\}_{i=1}^{L}\;\in\;\operatorname*{argmax}_{\{(\mu^{i},v^{i})\}_{i=1}^{L}}\mathbb{E}_{\pi_{\!\theta}}\bigl[\log\pi^{\theta}_{1}(X_{0})\bigr],$
where, due to the factorization of $\pi^{\theta}_{1}$, the loss writes equivalently as $\sum_{i=1}^{L}\mathbb{E}_{\pi^{i}_{\!\theta}}\bigl[\log\mathrm{N}(X_{0}^{i};\mu^{i}_{\theta},\mathrm{Diag}(v^{i}_{\theta}))\bigr].$ For each $i$, this is exactly the standard MLE problem for a multivariate Gaussian with diagonal covariance, whose one solution is given by matching the mean and per–coordinate variances of $\pi^{i}_{\!\theta}$; i.e. $\mu^{i}_{\theta}=\mathbb{E}_{\pi^{i}_{\!\theta}}[X^{i}_{0}]$ and $v^{i}_{\theta}=\mu^{i}_{\theta}\odot(1-\mu^{i}_{\theta})$. We restrict ourselves to a diagonal covariance in order to avoid expensive matrix inversions in the denoiser expression derived next. Data–dependent base distributions of this kind have also been considered in other applications, see for instance lee2022priorgrad; popov2021grad; ohayon2025posterior.
When using the base distribution $\pi^{\theta}_{1}$ and setting $\eta_{s}=0$ for all $s\in[0,1]$, the DDIM map (11) keeps the same form as before. The denoiser, however, is different and is now given in matrix form by $\hat{x}^{\theta}_{0}(x_{t},t)\!=\!\text{softmax}(\varphi_{\theta}+\frac{\alpha_{t}\lambda_{\theta}}{\sigma^{2}_{t}}\odot(x_{t}-\sigma_{t}\mu_{\theta}-\frac{\alpha_{t}}{2}\mathbf{1})).$ where $\lambda_{\theta}\in{\mathbb{R}}^{L\times K}$ with $\lambda^{i,j}_{\theta}=1/v^{i,j}_{\theta}$ and $\mathbf{1}\in{\mathbb{R}}^{L\times K}$ is the all-ones matrix. See Section C.2 for a derivation and Appendix C for the DDIM sampler with arbitrary schedule $(\eta_{s})_{s\in[0,1]}$. We refer to the resulting gradient estimator as ReDGE-Cov.
@@ 3.5 Related work
We discuss reparameterization tricks beyond the Gumbel-Softmax. potapczynski2020invertible replace Gumbel noise by Gaussian noise passed through an invertible transformation to obtain a more flexible family of continuous distributions on the simplex. wang2020relaxed go beyond the independence assumption (2) and propose a relaxation for correlated multivariate Bernoulli via a Gaussian copula. paulus2020gradient generalize the Gumbel–max trick by considering solutions to random linear programs and then obtain differentiable relaxations through by adding of strongly convex regularizer. Another way to obtain low variance gradient estimators is through the combination of REINFORCE with reparameterization trick-based control variates, or the use of Rao–Blackwellization (tucker2017rebar; grathwohl2018relax; liu2019rao; paulus2020raoblackwell).
Denoiser for mixture of Dirac delta.
When training a diffusion model using a dataset $(X_{i})_{i=1}^{N}$, the minimizer of the denoising loss is the denoiser for the empirical distribution $N^{-1}\!\sum_{i=1}^{N}\updelta_{X_{i}}$ and is available in closed form; see karras2022elucidating. Various recent works use this insight in different forms. scarvelis2023closed smooth the closed-form empirical denoiser to obtain training-free diffusion samplers that generalize beyond memorization. kamb2025an study denoising under architectural constraints,most notably equivariance and locality, and derive the optimal denoiser within this restricted function class, and show that the resulting training-free diffusion sampler generates novel samples and closely match the behavior of trained convolution-based diffusion models. ryzhakov2024explicit propose to train diffusion models by directly regressing to the empirical denoiser. In this work, we similarly leverage closed-form denoisers, here for distributions over $V^{L}$, but for a different goal: we use them to construct a soft reparameterization and differentiate through diffusion trajectories to obtain pathwise gradients. Finally, in concurrent work, andersson2025diffusion propose using diffusion models, in a sequential Monte Carlo setting, to generate $N$ i.i.d. reparameterized samples from the parameter-dependent empirical mixture $\smash{\sum_{i=1}^{N}w_{i}^{\theta}\,\delta_{X_{i}^{\theta}}}$, where $\smash{w_{i}^{\theta}\geq 0}$ and $\smash{\sum_{i=1}^{N}w_{i}^{\theta}=1}$, and $\theta$ denotes the state-space model parameters. This enables parameter estimation by differentiating end-to-end through the particle filter used to estimate the observation likelihood.

### v2

@@ 3.3 Extensions
We derive a ReinMax (Liu et al., 2023a) version of our diffusion-based reparameterization trick. First, by the marginalization property we have that $\pi_{\!\theta}(x_{0})=\int\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})\pi^{\theta}_{{t_{1}}}(x_{t_{1}})\mathrm{d}x_{t_{1}}$, and we can write, using the tower property, that $\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{E}[h_{\theta}(X_{t_{1}})]$, where $X_{t_{1}}$ is given by (7) with $\pi_{0}=\pi_{\!\theta}$ and $\smash{h_{\theta}(x_{t_{1}})\coloneqq\sum_{x_{0}}f(x_{0})\,\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})}$. The Gaussian mixture $\pi^{\theta}_{{t_{1}}}$ can be reparameterized approximately using the DDIM map $T^{\theta}_{t_{1}}$ in (10) and therefore $\mathbb{E}[h_{\theta}(X_{t_{1}})]\approx\mathbb{E}[h_{\theta}(T^{\theta}_{t_{1}}(X_{1}))]$ for any $\theta\in\Theta$. A Monte Carlo estimator of the total gradient of the r.h.s. at $\theta=\theta^{\prime}$ is given by
$\nabla_{\!{\theta}}h_{\theta}(T^{\theta^{\prime}}_{{t_{1}}}(X_{1})){\big|_{\theta=\theta^{\prime}}}+\operatorname{J}_{\theta}\!T^{\theta}_{t_{1}}(X_{1})^{\top}{\big|_{\theta=\theta^{\prime}}}\!\nabla_{\!{x}}h_{\theta^{\prime}}(T^{\theta^{\prime}}_{t_{1}}(X_{1})),$
@@
where the intractable terms are the gradients w.r.t. $\theta$ and $x$ of the conditional expectation $h_{\theta}$. The key observation is that for any $x_{t_{1}}$, the gradient of $\theta\mapsto h_{\theta}(x_{t_{1}})$ is a specific case of differentiating an expectation w.r.t. the parameters of a categorical distribution, which in this case is $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})$. Here by using the Straight-Through approximation (4) we recover our hard gradient estimator (ReDGE); i.e. $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})\approx\nabla_{\!{\theta}}f(\hat{x}^{\theta}_{0}(x_{t_{1}},{t_{1}}))$. Our ReinMax-based estimator replaces hard ST with ReinMax (46) as an estimator of $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})$. We refer to this gradient estimator as ReinDGE. When using a single diffusion step, i.e. ${t_{1}}=1$, the function $h_{\theta}$ is constant and equal to $\mathbb{E}_{\pi_{\!\theta}}[X]$ due to the boundary condition $\alpha_{1}=0$, and ReinMax is recovered as a special case. The same observation holds for the map $x_{t_{1}}\mapsto h_{\theta}(x_{t_{1}})$, for any $\theta$, but here we simply use the Straight-Through estimator.
Parameter dependent $\pi_{1}$.
In the previous construction, the terminal distribution $\pi_{1}$ is fixed to a standard Gaussian $\pi_{1}=\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$. In our setting, however, we can exploit the factorization (2) to select a parameter–dependent Gaussian distribution $\pi^{\theta}_{1}$ that best approximates $\pi_{\!\theta}$ in the maximum–likelihood sense. Specifically, we take $\pi^{\theta}_{1}$ with factorized density $\pi^{\theta}_{1}(x)=\prod_{i=1}^{L}\mathrm{N}(x^{i};\mu^{i}_{\theta},\mathrm{Diag}(v^{i}_{\theta}))$, where for all $i\in[L]$, $(\mu^{i}_{\theta},v^{i}_{\theta})\in{\mathbb{R}}^{K}\times{\mathbb{R}}^{K}_{>0}$ and $\mathrm{Diag}(v^{i}_{\theta})\in{\mathbb{R}}^{K\times K}$ is a diagonal matrix with $v^{i}_{\theta}$ as diagonal entries. The parameters are then defined as any solution to the maximum–likelihood problem of maximizing $\mathbb{E}_{\pi_{\!\theta}}[\log\pi^{\theta}_{\!1}(X_{0})]$ w.r.t. $(\mu_{\theta},v_{\theta})$ whose one solution is given by matching the mean and per–coordinate variances of $\pi^{i}_{\!\theta}$; i.e. $\mu^{i}_{\theta}=\mathbb{E}_{\pi^{i}_{\!\theta}}[X^{i}_{0}]$ and $v^{i}_{\theta}=\mu^{i}_{\theta}\odot(1-\mu^{i}_{\theta})$. We restrict ourselves to a diagonal covariance in order to avoid expensive matrix inversions in the denoiser expression derived next. Data–dependent base distributions of this kind have also been considered in other applications, see for instance Lee et al. (2022); Popov et al. (2021); Luo et al. (2023); Ohayon et al. (2025). When using the base distribution $\pi^{\theta}_{1}$ and setting $\eta_{s}=0$ for all $s\in[0,1]$, the DDIM map (9) keeps the same form as before. The denoiser, however, is different and is now given in matrix form by $\hat{x}^{\theta}_{0}(x_{t},t)\!=\!\text{softmax}(\varphi_{\theta}+\frac{\alpha_{t}\lambda_{\theta}}{\sigma^{2}_{t}}\odot(x_{t}-\sigma_{t}\mu_{\theta}-\frac{\alpha_{t}}{2}\mathbf{1})).$ where $\lambda_{\theta}\in{\mathbb{R}}^{L\times K}$ with $\lambda^{i,j}_{\theta}=1/v^{i,j}_{\theta}$ and $\mathbf{1}\in{\mathbb{R}}^{L\times K}$ is the all-ones matrix. See Section C.2 for a derivation and Appendix C for the DDIM sampler with arbitrary schedule $(\eta_{s})_{s\in[0,1]}$. We refer to the resulting gradient estimator as ReDGE-Cov. Finally, for large vocabularies $K$, the diagonal covariance can become ill-conditioned, so we also consider a scalar variant with $\mathrm{Diag}(v^{i}_{\theta})=\sigma_{\theta}^{2}\mathbf{I}_{K}$.
@@ 3.4 Related work
Beyond Gumbel-Softmax, several works propose alternative approximate reparameterizations. Potapczynski et al. (2020) replace Gumbel noise with an invertible push-forward of a Gaussian, yielding a richer family of simplex-valued relaxations. Wang and Yin (2020) relax the factorization assumption in (2) by modeling correlated multivariate Bernoulli variables via a Gaussian copula. Paulus et al. (2020b) generalize the Gumbel–max trick through solutions of random linear programs, obtaining differentiable relaxations by adding a strongly convex regularizer.
Denoiser for a mixture of Dirac deltas.
Dieleman et al. (2022) propose CDCD, which models categorical data by training a diffusion model on embedded tokens. Because the underlying variables are discrete, the denoiser is learned with a cross-entropy objective. When fitting a diffusion model to a finite dataset $(X_{i})_{i=1}^{N}$, the minimizer of the denoising objective is precisely the denoiser associated with the empirical distribution $N^{-1}\sum_{i=1}^{N}\updelta_{X_{i}}$, and it admits a closed-form expression; see Karras et al. (2022, Appendix B.3). Since our purpose is not to synthetise new data but to have differentiable sampling procedure, we similarly exploit closed-form denoisers for distributions on $\mathsf{V}^{L}$. In concurrent work, Andersson and Zhao (2025) propose using diffusion models within a sequential Monte Carlo setting to generate $N$ i.i.d. reparameterized samples from the parameter-dependent empirical mixture $\smash{\sum_{i=1}^{N}w_{i}^{\theta}\,\delta_{X_{i}^{\theta}}}$, where $\smash{w_{i}^{\theta}\geq 0}$ and $\smash{\sum_{i=1}^{N}w_{i}^{\theta}=1}$, and $\theta$ denotes the state-space model parameters. This enables parameter estimation by differentiating end-to-end through the particle filter used to estimate the observation likelihood.